Solve the differential equation by separation of variables.


(x^3 + 2)y = x(y^4 + 3)y’ ; at x = 1 and y = 1


1
Expert's answer
2021-09-17T03:45:22-0400

Let us solve the differential equation by separation of variables:

"(x^3 + 2)y = x(y^4 + 3)y'"; at "x = 1" and "y = 1". It follows that "(x^3 + 2)y = x(y^4 + 3)\frac{dy}{dx}," and hence "\frac{x^3 + 2}{x}dx = \frac{y^4 + 3}{y}dy." It follows that "\int\frac{x^3 + 2}{x}dx = \int\frac{y^4 + 3}{y}dy," and hence "\int(x^2+\frac{2}{x})dx = \int(y^3+\frac{3}{y})dy." Therefore, the general solution is "\frac{x^3}{3}+2\ln|x|=\frac{y^4}4+3\ln|y|+C." Since "y=1" for "x=1", we get that "\frac{1}{3}=\frac{1}4+C," and thus "C=\frac{1}{3}-\frac{1}4=\frac{1}{12}." We concludwe that the solution is "\frac{x^3}{3}+2\ln|x|=\frac{y^4}4+3\ln|y|+\frac{1}{12}."